An elastic collision of two pucks (mA = 0.5 kg, mg = 0.3 kg) on a frictionless air-jockey table. Puck A has an initial velocity of 4m/s in the positive X-direction and a final velocity of 2m/s in an unknown direction a. Puck B is initially at rest. Find the final speed of puck B and the angles a and b.


1
Expert's answer
2021-01-20T14:58:25-0500

For an elastic collision


X: "Mv=mu'\cdot \cos\beta+Mv'\cdot\cos\alpha"


Y: "mu'\cdot \sin\beta=Mv'\cdot\sin\alpha"


KE: "Mv^2/2=mu'^2/2+Mv'^2/2 \to"


"0.5\cdot4^2/2=0.3\cdot u'^2/2+0.5\cdot 2^2/2\to u'=4.47(m/s)" .Answer



"0.5\cdot 4=0.3\cdot 4.47\cdot \cos\beta+0.5\cdot 2\cdot\cos\alpha"


"0.3\cdot 4.47\cdot \sin\beta=0.5\cdot 2\cdot\sin\alpha\to" "\alpha\approx36.99°" and "\beta\approx26.66°" . Answer